OpenStudy (anonymous):

Which is the simplified form of the expression (P^2)(P^5)-4 times (P^-4)q^5)^-2 These are expressions ! Here are the chocies... A. 1/q^30 B. q^30 C. p^12/q^60 D. P^12 q^60

OpenStudy (anonymous):

\[((p ^{2})(p ^{5}-4)\times(p ^{-4}q ^{5})^{-2}\]

OpenStudy (anonymous):

Whats the answer?

OpenStudy (anonymous):

this will transformed into \[\frac{ (p ^{7}-4) }{ (\frac{ q ^{5} }{p ^{4} })^{2}}\]

OpenStudy (anonymous):

we don't give answers here, we hwlp students